OpenStudy (anonymous):

lim(x->infinity) ( sqrt(7+9x^2))/(1-2x). Am i supposed to rationalize the numerator?

OpenStudy (rulnick):

It's gonna be -3/2.

OpenStudy (rulnick):

In the limit, the 9x^2 dominates, so the numerator is effectively 3x.

OpenStudy (rulnick):

Likewise, denom -2x dominates.

OpenStudy (rulnick):

So 3/-2, or -3/2 is the lim.

OpenStudy (anonymous):

you basically ignore numbers that have little affect on the answer for example you ignore 7

OpenStudy (anonymous):

so all you have is sqrt ( 9x^2)/-2x

OpenStudy (anonymous):

3x/-2x = -3/2

OpenStudy (anonymous):

oh is that just a rule? the ration between the two factors with the biggest exponent?

OpenStudy (anonymous):

sound about right.

OpenStudy (anonymous):

well that explains a lot aha

OpenStudy (rulnick):

Be careful. You really have to look at all the terms. The art is in recognizing which terms will dominate close to the limit ... often pretty easy for "really big" x, when the lim is x -> infty.

OpenStudy (anonymous):

if it was goign to negative infinity would it change anything?

OpenStudy (rulnick):

Just the sign.

OpenStudy (anonymous):

the terms dominating would be the term with the biggest postive exponent, and the term with the biggest negative exponent?

OpenStudy (turingtest):

the any term with a negative x in the exponent will go to zero in the limit as x goes to positive infinity

OpenStudy (turingtest):

\[\lim_{x\to\infty}a^{-x}=0\]

OpenStudy (anonymous):

but when it goes to negative infinity?

OpenStudy (turingtest):

then the exponent approaches positive infinity, so the limit is positive infinity\[\lim_{x\to-\infty}a^{-x}=\infty\]

OpenStudy (anonymous):

oh! okay thanks

OpenStudy (turingtest):

welcome :D

OpenStudy (anonymous):

i feel like sometimes the way its taught is so confusing, when theres way easier ways to figure it out haha